# VideoGen-Eval: Agent-based System for Video Generation Evaluation

### v1

@@ 2.3 Measurements for Video Generation
Existing measurements can be categorized into two groups: objective metrics and human-aligned subjective metrics. Objective metrics typically assess a specific dimension via pre-trained priors. For instance, some methods utilize averaged frame-wise similarity metrics to calculate temporal consistency [48, 84, 10, 15, 68, 51]. Some leverage low-level motion analysis through optical flow [22, 14, 56, 58], frame interpolation [35], or tracking-based methods [55, 23] to measure the temporal variations. Some adopt vision-language models [34, 48, 33] to measure semantic alignment between generated content and multi-modal controls. Notwithstanding their merits, these measurements face challenges: 1) the smoothing effect in frame-level computations tends to obscure spatial-temporal localized anomalies; 2) inherent bias introduced by single-dimensional metrics. For instance, CLIP-score favors static scenes over legitimate motion patterns; 3) lack of explicit alignment with human perceptions. Human-aligned metrics typically construct datasets with human feedback, and then train quality assessment models based on them, such as video assessment models [70, 47, 26, 57, 71] and some image-oriented assessment models [24, 75, 74, 76, 59] that calculates the average of the frame-level subjective-aligned score. In this work, we utilize cutting-edge powerful generative models to build a high-quality dataset with human annotations and establish a multi-modal agent for comprehensive evaluation.
@@ 3 Evaluation Data Construction
Based on the above principles, we select a portion from the first part of 700 prompts and integrated prompts from multiple benchmarks (Vbench [18], MovieGen bench [45], T2V-CompBench [55]) to construct 700 content-rich prompts. Each prompt contains five key components: camera movement, background description, subject description, style, and lighting description. Note that some benchmark prompts do not cover all these components, so we use GPT-4o [1] to structurally expand the raw prompts, ensuring that each one contains these components. The word frequency in the final prompt is shown in Fig. 3. Finally, we use Sora [5], Hunyuan [20], Pixverse, Seaweed [7], Kling [27], Hailuo [42], Runway [50], and Wanx [61] to generate videos based on these prompts and conduct the benchmark.
@@ 4.3 Temporal Patch Tools
Figure 6: a The distribution of scores given by humans and agent systems to each dimension of 8 models (The top denotes the human, and the bottom denotes the agent system). As well as the specific scores, (b) denotes the human, and (c) denotes the agent.
@@ 5.3 Image-to-Video Evaluation
Wanx Kling Hailuo Pixvers. Seawe. Sora Runway
@@ 6 Conclusion
We introduce VideoGen-Eval, a novel agent-based dynamic evaluation framework that integrates structured content, MLLM-based judgment, and patch tools to assist in addressing the gap between recent video generation and evaluation. Coupled with our benchmark, this evaluation system exhibits comprehensive and human-aligned evaluation of cutting-edge video generation models. This establishes a dynamically extensible paradigm for video generation evaluation. As MLLMs and video generation models continue to co-evolve, the proposed system framework enables more robust evaluation capabilities while progressively aligning with human preferences.

### v2

@@ 2.3 Measurements for Video Generation
Existing measurements can be categorized into two groups: objective metrics and human-aligned subjective metrics. Objective metrics typically assess a specific dimension via pre-trained priors. For instance, some methods utilize averaged frame-wise similarity metrics to calculate temporal consistency [48, 85, 10, 15, 68, 51]. Some leverage low-level motion analysis through optical flow [22, 14, 56, 58], frame interpolation [35], or tracking-based methods [55, 23] to measure the temporal variations. Some adopt vision-language models [34, 48, 33] to measure semantic alignment between generated content and multi-modal controls. Notwithstanding their merits, these measurements face challenges: 1) the smoothing effect in frame-level computations tends to obscure spatial-temporal localized anomalies; 2) inherent bias introduced by single-dimensional metrics. For instance, CLIP-score favors static scenes over legitimate motion patterns; 3) lack of explicit alignment with human perceptions. Human-aligned metrics typically construct datasets with human feedback, and then train quality assessment models based on them, such as video assessment models [70, 47, 26, 57, 71] and some image-oriented assessment models [24, 75, 74, 76, 59] that calculate the average of the frame-level subjective-aligned score. The Evaluation Agent [84] is most similar to ours, which integrates the proposal and execution stages using an agent and completes the evaluation. In this work, we utilize cutting-edge powerful generative models to build a high-quality dataset with human annotations and establish a multi-modal agent for comprehensive evaluation.
@@ 3 Evaluation Data Construction
Based on the above principles, we select a portion from the first part of 700 prompts and integrated prompts from multiple benchmarks (Vbench [18], MovieGen bench [45], T2V-CompBench [55]) to construct 700 content-rich prompts. Each prompt contains five key components: camera movement, background description, subject description, style, and lighting description. Note that some benchmark prompts do not cover all these components, so we use GPT-4o [1] to structurally expand the raw prompts, ensuring that each one contains these components. The word frequency in the final prompt is shown in Fig. 3. Finally, we use Sora [5], Hunyuan [20], Pixverse, Seaweed [7], Kling [27], Hailuo [42], Runway [50], and Wan2.1 [61] to generate videos based on these prompts and conduct the benchmark.
@@ 4.3 Temporal Patch Tools
Figure 6: (a) The distribution of scores given by humans and agent systems to each dimension of 8 models (The top denotes the human, and the bottom denotes the agent system). As well as the specific scores, (b) denotes the human, and (c) denotes the agent.
@@ 5.3 Image-to-Video Evaluation
Wan2.1 Kling Hailuo Pixvers. Seawe. Sora Runway
@@ 6 Conclusion
We introduce VideoGen-Eval, a novel agent-based dynamic evaluation framework that integrates structured content, MLLM-based judgment, and patch tools to assist in addressing the gap between recent video generation and evaluation. Coupled with our benchmark, this evaluation system exhibits a comprehensive and human-aligned evaluation of cutting-edge video generation models. This establishes a dynamically extensible paradigm for video generation evaluation. As MLLMs and video generation models continue to co-evolve, the proposed system framework enables more robust evaluation capabilities while progressively aligning with human preferences.